# Non-linear anomalous Edelstein response at altermagnetic interfaces

### v2

@@ Appendix D Boltzmann approach to non-linear EE
Edelstein response up to second order is given by
$\expectationvalue{\sigma_{k}}=\chi_{ik}E_{i}+\tilde{\chi}_{ijk}E_{i}E_{j},$ (13)
where $\expectationvalue{\sigma_{k}}$ is the magnetization in the $k$ direction, $E_{i}$ is the $i-$th component of the electric field, and $\chi_{ij}$ is the linear Edelstein susceptibility, while $\tilde{\chi}_{ijk}$ is the second order susceptibility. We can also define the magnetization $\expectationvalue{\sigma_{k}}$ also given by
$\expectationvalue{\sigma_{k}}=\mu_{b}\sum_{n\hskip 1.42271ptocc}\int\frac{d^{2}\mathbf{k}}{(2\pi)^{2}}s^{n}_{k}(\mathbf{k})f(\mathbf{k}),$ (14)
where $s_{k}^{n}$ is the $k$-th spin component for the $n$-th band and $f(\mathbf{k})$ is the complete thermal distribution.
In the Boltzmann approach, the equation for the particle distribution $f$ is the following
$\frac{\partial f}{\partial t}+\frac{d\mathbf{r}}{dt}\cdot\frac{\partial f}{\partial\mathbf{r}}+\frac{d\mathbf{k}}{dt}\cdot\frac{\partial f}{\partial\mathbf{k}}=\text{Coll},$ (15)
where Coll$=-\frac{\varphi}{\tau}$, with $\tau$ being the relaxation time that we are considering a constant, $\varphi=f-f_{th}$, and $f_{th}$ the Fermi-Dirac distribution. To first order, for stationary transport and in homogeneous materials, the previous equation becomes
$-|e|\mathbf{E}\cdot{\nabla}_{\mathbf{k}}f_{th}=-\frac{\varphi^{(1)}}{\tau},$ (16)
where $\varphi^{1}$ is the correction to the particle distribution at the first order and $e$ is the electron charge. Solving Eq. (16) leads to
$\varphi^{(1)}=|e|\tau\mathbf{E}\cdot\mathbf{\nabla}_{\mathbf{k}}f_{th}.$ (17)
For the next order in the electric field, it is necessary to start again from the Boltzmann equation, which now takes the form
$-|e|\mathbf{E}\cdot{\nabla}_{\mathbf{k}}\varphi^{(1)}=-\frac{\varphi^{(2)}}{\tau},$ (18)
where $\varphi^{(2)}$ is the the correction to the particle distribution at the second order in the electric field $\mathbf{E}$. We neglect the effect of the magnetic field $\mathbf{B}$ since, by considering an in-plane magnetic field in a 2D system, we have $(\mathbf{v}\times\mathbf{B})\cdot\nabla_{\mathbf{k}}\varphi^{(1)}=0$.
Figure 14: Second-order Edelstein susceptibility $\tilde{\chi}_{xxz}$ as a function of $\alpha$ for $\mu=0$ and $\gamma=1.881$ eV Å2. The line is a guide to the eye.
Substituting Eq. (17) into Eq. (18), and considering a constant $\tau$, $\varphi^{(2)}$ can be written as
$\varphi^{(2)}=e^{2}\tau^{2}\frac{\partial}{\partial k_{j}}\left(\frac{\partial f_{th}}{\partial k_{i}}\right)E_{i}E_{j}.$ (19)
Therefore, we obtain
@@
We can estimate the coefficient $\tilde{\chi}_{0}=\frac{\mu_{B}e^{2}\tau^{2}}{4\pi^{2}}S_{\text{cell}}$, with $a$ the characteristic length of our system, which will be the lattice parameter, $S_{\text{cell}}$ to be the area of the unit cell, as
$\tilde{\chi}_{0}=2.36\times 10^{-14}\;\frac{\text{m}^{2}}{\text{V}^{2}}.$ (21)
It is straightforward to demonstrate the necessity of the RSOC in order to have a non-vanishing $\tilde{\chi}_{xxz}$. We can consider the integrand of Eq. (20) without the derivative of the Fermi function as
$G^{\pm}(k_{x},k_{y})=\pm v_{x}^{\pm}(k_{x},k_{y})\frac{\partial s_{z}(k_{x},k_{y})}{\partial k_{x}}.$ (22)
The group velocity (corresponding to the system of Eq. (1) for $\alpha=0$) is
$v_{x}^{\pm}(k_{x},k_{y})=\frac{k_{x}}{m}\pm 2\gamma k_{x}\mathrm{sgn}{(k_{x}^{2}-k_{y}^{2})},$ (23)
while the spin for the two bands can be written as
$s_{z}(k_{x},k_{y})=\mathrm{sgn}(k_{x}^{2}-k_{y}^{2}),$ (24)
which corresponds to the following derivative
$\frac{\partial s_{z}(k_{x},k_{y})}{\partial k_{x}}=4k_{x}\delta(k_{x}^{2}-k_{y}^{2}).$ (25)
The integrand for the two bands is therefore {dmath} G^±(k_x,k_y)=±4 k_x^2 δ(k_x^2-k_y^2) + 8γk_x^2 δ(k_x^2-k_y^2)sgn(k_x^2-k_y^2). The presence of $\delta(k_{x}^{2}-k_{y}^{2})$ makes this integrand non-zero only along the $k_{x}=\pm k_{y}$ lines; however, since the full integrand is ±$\sum_{\pm}G^{\pm}(k_{x},k_{y})\delta(\varepsilon^{\pm}(k_{x},k_{y})-\mu)$, at fixed $\mu$ along these lines only the four points of degeneracy of the two bands contribute to the conductance. Moreover, since they are degenerate, we can write the sum of $G^{+}$ and $G^{-}$ as {dmath} G=G^+(k_x,±k_x)+G^-(k_x,±k_x)=16 γk_x^2 sgn(k_x^2-k_x^2). However, the sign function $(\mathrm{sgn})$ on these lines corresponds to the exact point where the function has a discontinuity. In this case, we can either set $\mathrm{sgn}(0)=0$, or consider the sum of the left and right limits of the $\mathrm{sgn}$ function. In either case, the sum of these points leads to a vanishing the conductance.
Appendix E Temperature dependence
Figure 15: Comparison of the calculated Edelstein susceptibility for $T=33$ K and $T=328$ K, considering the parameters $\gamma=1.881$ eV Å2, $\alpha=52$ meV Å, and $h_{y}=40$ eV.
For higher temperatures, Figure 15 displays $\chi_{xi}(\mu)$ for two different benchmark values. Increasing the temperature from $33$ K to $328$ K preserves the overall behavior of the susceptibility by changing the chemical potential reported in Fig. 4(c), only smoothing out the curves, especially at low fillings.

### v3

@@ Appendix D Boltzmann approach to non-linear EE
The first- and second-order Edelstein reads
$M_{k}=\chi_{ik}E_{i}+\tilde{\chi}_{ijk}E_{i}E_{j},$ (10)
where $\chi_{ik}$ and $\tilde{\chi}_{ijk}$ are the linear and second-order Edelstein susceptibilities, respectively. In the Boltzmann approach, the magnetization is obtained as
$M_{k}=\mu_{B}\sum_{n\hskip 1.42271ptocc}\int\frac{d^{2}\mathbf{k}}{(2\pi)^{2}}s^{n}_{k}(\mathbf{k})f(\mathbf{k}),$ (11)
with $s_{k}^{n}$ the spin component of band $n$ and $f(\mathbf{k})$ the distribution function.
In the Boltzmann framework, we expand $f=f_{\rm th}+\varphi^{(1)}+\varphi^{(2)}+\cdots$ in powers of $\mathbf{E}$, with the lowest-order $f_{\rm th}$ being the Fermi distribution function, and for homogeneous, stationary transport with constant relaxation time $\tau$, the first- and second-order corrections read
$\varphi^{(1)}$ $=|e|\tau\,\mathbf{E}\cdot\nabla_{\mathbf{k}}f_{\rm th},$ (12)
$\varphi^{(2)}$ $=e^{2}\tau^{2}\,\frac{\partial}{\partial k_{j}}\left(\frac{\partial f_{\rm th}}{\partial k_{i}}\right)E_{i}E_{j}.$ (13)
The second-order susceptibility is then given by
@@
An estimate of the prefactor yields
$\tilde{\chi}_{0}=\frac{\mu_{B}e^{2}\tau^{2}}{4\pi^{2}}S_{\text{cell}}\simeq 2.36\times 10^{-14}\;\frac{\text{m}^{2}}{\text{V}^{2}}.$ (15)
We adopt the rigid band approximation, which is adequate for the purpose of our model. In principle, for a more complex multiband structure, corrections to the eigenstates due to a strong (or oscillating) electric field could be considered [26, 23], but these effects are not critical in our approach.
Figure 9: (a) Band structure of the tight-binding model of Hamiltonian (16) using: $t=0.6$, $\gamma=0.2$, $t_{1}=0.05$, and $\alpha=0.01$. The energy has been shifted in order to set the minimum of the bands to $\mu=0$. (b) Edelstein susceptibility $\chi_{xy}/\chi_{0}$ as a function of the chemical potential.
In the absence of Rashba coupling it is straightforward to see that $\frac{\partial s_{z}}{\partial k_{x}}=0$ and therefore $\chi_{xxz}=0$. We use these expressions to determine the non-linear Edelstein response shown in the main text.
Appendix E Tight-binding model
In this section we highlight the role of the crossing at high-symmetry points and the multi-band crossings on the linear Edelstein response. We construct a minimal tight-binding model from the irreducible representations of the altermagnetic symmetry group adopted in the main text, including the inversion-breaking Rashba term
$H(\mathbf{k})=[-2t(\cos k_{x}+\cos k_{y})\,-\mu]\sigma_{0}\tau_{0}-2\gamma(\cos k_{x}-\cos k_{y})\,\sigma_{z}\tau_{x}-2t_{1}(\cos k_{x}+\cos k_{y})\,\sigma_{0}\tau_{x}+\alpha\left[\sin k_{x}\,\sigma_{y}-\sin k_{y}\,\sigma_{x}\right]\tau_{x}.$ (16)
Here $t_{1}$ has been included to split the four-fold degeneracy. For $t_{1}=0$ and $t=1/2m$ the the expansion at ${\mathbf{k}}\approx 0$ is equivalent to Ham. (6). The band structure is shown in Fig. 9, together with the Edelstein response function for varying filling of the bands. The dominant contribution to the Edelstein response comes from the band crossings, as one could expect. In particular, the band crossings at the $\Gamma$ points, which give the largest contributions, are well described by our $\bf k\cdot p$ local models, providing a concrete example of the local nature of the Edelstein response. At large fillings, a specular contribution comes from the band crossings at X and Y, due to a change of chirality of the spin texture close to the corner of the Brillouin zone. Band crossings at M and M’ do not produce an Edelstein response, since the Rashba operator has vanishing matrix elements between the doubly degenerate eigenstates. This is immediately visible by expanding the Hamiltonian (16) $\mathbf{q}_{\Delta\mathbf{k}}={\bf k}-\Delta{\bf k}$ with $\Delta{\bf k}=(\pm\pi,0)$
$H(\mathbf{q}_{(\pi,0)})\approx t(-q_{x}^{2}+q_{y}^{2})\sigma_{0}\tau_{0}+\gamma(4-q_{x}^{2}-q_{y}^{2})\sigma_{z}\tau_{x}+t_{1}(-q_{x}^{2}+q_{y}^{2})\sigma_{0}\tau_{x}+\alpha(-q_{x}\sigma_{y}-q_{y}\sigma_{x})\tau_{x},$ (17)
so for $\mathbf{q}\to 0$ the doubly degenerate eigenstates with $\sigma_{z}\tau_{x}=+1$ (or $\sigma_{z}\tau_{x}=-1$) lead to a zero matrix element for the RSOC.
Another band crossing occurs at $k_{x}=k_{y}=\pi/2$; a local expansion gives
$H(\mathbf{q}_{(\pi/2,\pi/2)})\approx 2t(q_{x}+q_{y})\sigma_{0}\tau_{0}+2\gamma(q_{x}-q_{y})\sigma_{z}\tau_{x}+2t_{1}(q_{x}+q_{y})\sigma_{0}\tau_{x}+\alpha((1-q_{x}^{2})\sigma_{y}+(-1+q_{y}^{2})\sigma_{x})\tau_{x}.$ (18)
For $q\to 0$, the Rashba term reduces to $\alpha(\sigma_{y}-\sigma_{x})\tau_{x}$, so the crossing is gapped by the RSOC which enforces an in-plane spin. At finite $\bf q$, the spin tilts out of plane due to the altermagnetic terms, enforcing a spin-momentum locking and inducing a symmetry-allowed Edelstein response, which is clearly visible at half-filling, albeit of lower intensity than the response coming from the $\Gamma$ and X points.
Appendix F Intrinsic interband contribution
In this section, we discuss the role of the interband intrinsic contribution to the Edelstein effect in our model. Its general definition [32] is
$\chi^{\text{int}}_{ij}=e\hbar\int\frac{d^{2}{\bf k}}{(2\pi)^{2}}\sum_{m\neq n}\textrm{Im}\left[\frac{s_{j}^{nm}(\mathbf{k})v_{i}^{mn}(\mathbf{k})}{(\varepsilon_{\bf k}^{n}-\varepsilon_{\bf k}^{m})^{2}}\right]\bigl[f(\varepsilon_{\bf k}^{n})-f(\varepsilon_{\bf k}^{m})\bigr],$ (19)
[…]
This contribution has to be compared with the extrinsic Edelstein susceptibility
$\chi^{\mathrm{ext}}_{ij}=e\tau\sum_{n}\int_{\mathrm{BZ}}\frac{d^{2}{\bf k}}{(2\pi)^{2}}\;\Big(-\frac{\partial f_{0}}{\partial\varepsilon^{n}_{\mathbf{k}}}\Big)\,v^{n}_{i}(\mathbf{k})\,s^{n}_{j}(\mathbf{k}),$ (20)
where all quantities are expressed in standard units. From these definitions, it becomes immediately clear that the intrinsic contribution can dominate over the extrinsic one only over those regions of the Brillouin zone where $\varepsilon_{m}-\varepsilon_{n}\lesssim\hbar/\tau$ (excepting cases where the extrinsic contribution vanishes due to some symmetry, which we have established not to be the case for the components of the Edelstein susceptibility we determine). Since $\hbar/\tau\sim 0.7\,\mathrm{meV}$, whereas all the energy scales within our model Hamiltonian are significantly larger, clearly the interband intrinsic contribution can become competitive only close to near-degeneracy points. Let us therefore consider such contribution. If Zeeman terms are included, the spectrum is fully gapped, with typical gaps larger than $\sim 50\,\mathrm{meV}$. In this regime, the intrinsic susceptibility is always negligible compared to the extrinsic one. The only case requiring more careful analysis is the gapless limit at ${\bf k}=0$. By considering Hamiltonian (1), and using Eq. (19), one obtains for the integrand of $\chi_{ix}=e\hbar\int I(s_{i})\,d^{2}{\bf k}$
$I(s_{z})=-\frac{1}{4\pi^{2}}\frac{k_{y}\alpha^{2}}{4\left[\bigl(k_{x}^{2}+k_{y}^{2}\bigr)\alpha^{2}+\bigl(k_{x}^{2}-k_{y}^{2}\bigr)^{2}\gamma^{2}\right]^{3/2}},$ (21)
and
$I(s_{y})=k_{x}k_{y}\,\mathcal{F}(k_{x}^{2},k_{y}^{2}),$ (22)
where $\mathcal{F}(k_{x}^{2},k_{y}^{2})$ is an even function of $\bf k$}. Both integrands vanish upon Brillouin-zone integration by symmetry, so the intrinsic contribution is absent in this two-band model.